# Resonant spin Hall effect in a nano-ribbon of a spin-orbit coupled electron system

### v1

@@ IV.1 Spin Hall conductivity
Next, we briefly discuss the spin Hall phenomena in a $2D$ electronic system with an anisotropic RSOI and DSOI. The ab-initio calculation based theoretical prediction in the surface states of $Au(110)$ suggested the existence of an anisotropic Rashba splitting [39]. In fact the $C_{2\nu}$ point-group symmetry analysis not only suggested the mass anisotropy but also hinted an anisotropic RSOI [40] as $H_{R}=\alpha_{x}k_{x}\sigma_{y}+\alpha_{y}k_{y}\sigma_{y}$ where $\{\alpha_{x},\alpha_{y}\}$ are two independent parameters describing the degree of spin splitting along two orthogonal directions in momentum space. Similarly, anisotropic Dresselhaus interaction can be modeled as $H_{D}=\beta_{x}k_{x}\sigma_{x}-\beta_{y}k_{y}\sigma_{y}$ with $\{\beta_{x},\beta_{y}\}$ denoting the anisotropy in the DSOI. These generalization naturally extend the isotropic case and allow us to investigate the robustness of SHR under reduced lattice symmetry. In the discrete lattice model, such anisotropy can be captured by considering a weak imbalance between hoping parameters along two orthogonal directions as $t_{x}$ and $t_{y}$, as $t_{y}=\nu t_{x}$ where $\nu$ describes the degree of anisotropy. Following the same approach as in the isotropic Rashba–Dresselhaus case, we compute both the band structure and the corresponding SHC, and note that the spin degeneracy points and the anti-crossing points persist even in the presence of anisotropy. The SHC continues to exhibit resonance whenever the chemical potential lies at these points. It indicates that inter sub-bands resonance phenomena in spin Hall conductivity is robust to weak anisotropy in band structure.
@@
We now turn to the role of finite temperature on the SHC, we find that the suppression of spin Hall conductivity at finite temperature is more pronounced at the lower energy degeneracy points compared to the higher energy ones [Fig.4].This behavior can be understood in terms of the difference between the Fermi occupation factors. When the chemical potential lies in the lower bands, the upper band remains nearly unoccupied and hence the difference is essentially governed by the lower band Fermi function. In this regime, the thermal broadening of the Fermi function strongly reduces the SHC resonance. However, when the chemical potential is located in the higher bands, both the upper and lower bands are partially occupied. The effect of broadening then tends to cancel out between the two bands, leading to a much weaker suppression of SHC at the higher energy degeneracy and anticrossing points. Hence, we conclude that the resonance features in the higher bands are more robust to temperature effects.
@@
Next, we briefly check whether the longitudinal transport exhibits any signatures of vanishing spin splitting or anti-crossings.
IV.2 Longitudinal transport
@@
The longitudinal conductivity characterizes the charge transport along the translationally invariant direction under the application of an external bias. In order to capture the contribution of finite width induced sub-bands, it is convenient to use the real space Green function based approach developed by Sancho et al. [41, 42]. In this approach the longitudinal direction is taken to be very long compared to the width, and both the ends of the ribbon are attached to two leads of identical system. Following this approach it is straight forward to write the transmission between two leads though central region at $E$ as[43]
@@
At zero temperature, the longitudinal conductance follows from the Landauer–Büttiker relation[44, 45],
@@
Figure 5: Band structure (left) and corresponding longitudinal conductance (right, in units of $e^{2}/h$ ) for a zigzag edge nanoribbon with $\alpha=0.1$, $\beta=0.09$ in units of tata and the width is $N=21$. The brown dashed lines highlight interband anti-crossings while the black dashed lines show interband spin degeneracy point.
By using Eq. (12) we compute zero temperature conductance as a function of chemical potential against the corresponding band structure for zigzag edge nanoribbon in the Fig. 5. As usual, the conductance exhibits quantization in units of $2e^{2}/h$ when $\mu<0.1$ . This can be attributed to the fact that only one sub-band with two electron incident channels is occupied by the chemical potential. As the chemical potential starts occupying the spin up sub-band of the lowest sub-band, both bands contribute one incident channel each maintaining the conductance again at $2e^{2}/h$. At $\mu\approx 0.1$ the conductance suddenly increases to $4e^{2}/h$ reflecting the existence of four electron incident channels, three from the spin up sub-band and one from the spin down sub-band. However, within very a very narrow range of $\mu$ around $0.1$, the conductance drops back to $2e^{2}/h$ and then rises again to $4e^{2}/h$. This unusual feature originates from the presence of an anti-crossing point at $\mu=0.1$. With further increase in $\mu$, the phenomenon reappears at $\mu=0.3$ corresponding to another anticrossing point. We emphasize that such behavior does not occur at the inter-sub-band spin degeneracy points. Hence we can conclude that although a signature of anti-crossing point is visible in longitudinal conductance, it shows no corresponding features at the inter-band spin degeneracy. Finally, we note that the qualitative features of the conductance remain the same for the straight edge case, with the only difference being a shift in the locations of the anticrossing points.
@@ V Summary
We have investigated the spin-Hall transport and longitudinal transport for a 2D electronic system with RSOI and DSOI in a discrete lattice model. We have considered square lattice system with straight edge and zigzag edge geometry. In both cases, we showed that a number of spin degeneracy points and anti-crossing points between two nearest sub-band with opposite spin emerges. The spin-Hall conductivity shows a sharp enhancement when the chemical potential is situated at those spin-degeneracy points or anticrossing points, exhibiting spin-Hall resonance. Hence, we conclude that resonance in spin-Hall conductivity can be achieved without applying any external perturbation like magnetic field or light field as proposed previously. However such resonance can occur only if RSOI and DSOI both are present. We have also investigated the effects of anisotropy and finite temperature and demonstrated that the latter results in suppression of SHC but this suppression is weaker in higher energy degeneracy and anticrossing points while the former results in a shift of the position of resonance points in the spin Hall conductivity and finally we study the longitudinal conductance by using retarded Green function approach, and reveal that anti-crossing point can give a signature to the conductance whereas the degeneracy point does not leave any mark.
@@ VI Acknowledgement
SK Firoz Islam acknowledge financial support by the project ANRF/ECRG/2024/005166/PMS.
Appendix A Straight-edge nanoribbon: Band structure and Transport properties
We briefly discuss the case of a straight edge nanoribbon here. We present lowest few spin-split sub-bands for straight edge nanoribbon in Fig.6 where red and blue curves represent the spin-up and spin-down components, respectively. Very much similar to the case of zigzag ribbon, here also a number of spin-splitting vanishing as well as anti-crossing points are present between two nearest sub-bands, which are highlighted by circles and squares respectively. It is noticeable that the number of such points are relatively higher in straight edge case compared to zigzag edge. It is expected that the SHR can occur when the chemical potential passes through the inter sub-band spin degenerate or anti-crossing points. The SHC is plotted in the Fig. 7 showing multiple resonances.
@@ Appendix A Straight-edge nanoribbon: Band structure and Transport properties
Figure 6: Band dispersion for straight edge nanoribbon along $k_{y}$ for $\alpha=0.11$ and $\beta=0.1$ in units of tata. Red and blue curves denote spin-up and spin-down components. The emergence of interband anticrossings and degeneracy induced by the competition between Rashba and Dresselhaus couplings are highlighted by squares and circles respectively.
@@
Figure 7: Band structure (left) and corresponding spin Hall conductivity normalized with $\sigma_{0}=e/8\pi$ for straight edge nanoribbon with width $N=21$ and $\alpha=0.11$, $\beta=0.1$ in units of tata and the temperature $k_{B}T/t=10^{-5}$. The SHC exhibits resonance at various $\mu$. The brown dashed lines highlight the resonance due to interband anti-crossings, while the black dashed lines indicate the resonance due to vanishing spin splitting.
@@
Figure 8: Band structure (left) and corresponding longitudinal conductance (right, in units of $e^{2}/h$ ) for a straight edge nanoribbon with $\alpha=0.11$, $\beta=0.1$ in units of tata and the width is $N=21$. The brown dashed lines highlight interband anti-crossings while the black dashed lines show interband spin degeneracy point.

### v2

@@ IV.1 Spin Hall conductance
Next, we briefly discuss the spin Hall phenomenon in a 22D electronic system with an anisotropic RSOI and DSOI. The ab-initio calculation based theoretical prediction in the surface states of $Au(110)$ suggested the existence of an anisotropic Rashba splitting Nagano et al. (2009). In fact the $C_{2\nu}$ point-group symmetry analysis not only suggested the mass anisotropy but also hinted an anisotropic RSOI Nuber et al. (2008) as $H_{R}=\alpha_{x}k_{x}\sigma_{y}+\alpha_{y}k_{y}\sigma_{y}$ where $\{\alpha_{x},\alpha_{y}\}$ are two independent parameters describing the degree of spin splitting along two orthogonal directions in momentum space. Similarly, anisotropic Dresselhaus interaction can be modeled as $H_{D}=\beta_{x}k_{x}\sigma_{x}-\beta_{y}k_{y}\sigma_{y}$ with $\{\beta_{x},\beta_{y}\}$ denoting the anisotropy in the DSOI. In the discrete lattice model, such anisotropy can be captured by considering a weak imbalance between hopping parameters along two orthogonal directions as $t_{x}$ and $t_{y}$, as $t_{y}=\nu t_{x}$ where $\nu$ describes the degree of anisotropy. Following the same approach as in the isotropic Rashba–Dresselhaus case, we note that the spin degeneracy points and the anticrossing points persist even in the presence of anisotropy. The SHC continues to exhibit resonance whenever the chemical potential lies at these points. It indicates that inter-subband resonance phenomenon in SHC is robust to weak anisotropy in band structure.
@@
We now turn to the role of finite temperature on the SHC, we find that the suppression of SHC at finite temperature is more pronounced at the lower energy degeneracy points compared to the higher energy ones Fig. 4. This behavior can be understood in terms of the difference between the Fermi occupation factors. When the chemical potential lies in the lower bands, the upper band remains nearly unoccupied and hence the difference is essentially governed by the lower band Fermi function. In this regime, the thermal broadening of the Fermi function strongly reduces the SHC resonance. However, when the chemical potential is located in the higher bands, both the upper and lower bands are partially occupied. The effect of broadening then tends to cancel out between the two bands, leading to a much weaker suppression of SHC at the higher energy degeneracy and anticrossing points. Hence, we conclude that the resonance features in the higher bands are more robust to temperature effects.
We comment here that the typical sample that could be suitable for resonance phenomenon is InAs quantum wells in which both RSOI and DSOI with almost the same order of strength has been already confirmed by photocurrent measurement Ganichev et al. (2004). What we require is to reduce the width to engineer a number of subbands. To achieve the resonance in SHE, it is important to perform the fine tuning of chemical potential by gate voltage so that it capture the inter subband spin degenerate or anticrossing points. The experimentally relevant nanoribbon widths typically lie in the range of 50150nm50–150nm Scherübl et al. (2016). For InAs square lattice discretization with lattice constant $a=0.6nm$, a ribbon width of $W=77nm$ corresponds to $N\approx 128$ lattice sites across the transverse direction, placing our simulations well within experimentally accessible regimes. The strength of RSOI in InAs is known to be relatively large and electrically tunable, with experimentally reported values in the range $\alpha=(0.5-3)\times 10^{-11}eVm$, and can be tuned by gating. Using the InAs effective mass $m^{*}=0.02m_{e}$ Sladek (1957) and lattice spacing of InAs, experimentally reported Rashba couplings correspond to a dimensionless lattice Rashba strength $\alpha/t=10^{-3}-10^{-5}$. Importantly, while a single global gate generally changes both the chemical potential $\mu$ and the interfacial electric field and hence $\alpha$ experiments employing dual gate geometries have demonstrated that these two effects can be largely decoupled Scherübl et al. (2016); Fan et al. (2022); Liang and Gao (2012). In realistic device geometries, $\alpha$ can be varied independently of $\mu$, justifying our theoretical approach of treating the chemical potential and Rashba coupling as independently tunable parameters.
IV.2 Spin Nernst coefficient
Having evaluated the SHC of the nanoribbon, we now extend our analysis to its thermal analogue by computing the spin Nernst coefficient. The SNC characterizes the transverse spin current generated in response to a longitudinal temperature gradient, and is defined as $\alpha^{z}_{xy}=-J^{z}_{x}/\nabla_{y}T$. Within linear response theory, SNC is obtained as Guo and Wang (2017); Fujimoto et al. (2024); Wimmer et al. (2013)
$\alpha^{z}_{xy}=\frac{1}{T}\sum_{\zeta}\Omega^{z,\zeta}_{xy}\Lambda_{\zeta}$ (12)
with
$\Omega^{z,\zeta}_{xy}$ $=-2\,\mathrm{Im}\sum_{\zeta^{\prime}\neq\zeta}\frac{\langle\zeta|\hat{J}^{z}_{x}|\zeta^{\prime}\rangle\langle\zeta^{\prime}|\hat{v}_{y}|\zeta\rangle}{\left(E_{\zeta}-E_{\zeta^{\prime}}\right)^{2}},$ (13)
$\Lambda_{\zeta}$ $=(E_{\zeta}-\mu)\,f_{\zeta}+k_{B}T\,\ln\!\bigl(1+e^{-\beta(E_{\zeta}-\mu)}\bigr)$
Figure 5: spin resolved band structure (left) and spin Nernst coefficient normalized with $\alpha_{xy}^{z,0}=-k_{B}/(4\pi)$ for a zigzag edge nanoribbon with width $N=21$ and $\alpha=0.1$, $\beta=0.09$ in units of tata at temperature $k_{B}T/t=0.01$.
The computed SNC as a function of the chemical potential, together with the corresponding band structure, is shown in Fig. 5. We observe a sequence of sharp positive and negative peaks that occur precisely when the chemical potential crosses the anticrossing or spin degeneracy points in the spectrum. These features are fully consistent with the behaviour of the SHC, which exhibits spin Hall resonance at the same points.
$\alpha^{z}_{xy}=-\,\frac{\pi^{2}}{3e}\,k_{B}^{2}T\,\sigma^{z}_{xy}(\mu)^{\prime}$ (14)
This behavior is also consistent with the thermoelectric Mott relation, Eq. (14)  Guo and Wang (2017), which links the SNC to the energy derivative of the SHC. Thus, whenever the SHC undergoes a rapid change such as at anticrossings or near spin degenerate points, its derivative becomes large, giving rise to the pronounced peaks in the SNC. The overall structure of the SNC therefore provides a clear thermoelectric fingerprint of the underlying anticrossing and spin degenerate points.
@@ IV.2 Spin Nernst coefficient
Next, we briefly check whether the longitudinal transport exhibits any signatures of vanishing spin splitting or anticrossings.
IV.3 Longitudinal transport
@@
The longitudinal conductivity characterizes the charge transport along the translationally invariant direction under the application of an external bias. In order to capture the contribution of finite width induced subbands, it is convenient to use the real space Green function based approach developed by Sancho et al. Sancho et al. (1984, 1985). In this approach the longitudinal direction is taken to be very long compared to the width, and both the ends of the ribbon are attached to two leads of identical system. Following this approach it is straight forward to write the transmission between two leads though central region at $E$ as Caroli et al. (1971)
@@
At zero temperature, the longitudinal conductance follows from the Landauer–Büttiker relation (Landauer, 1970; Büttiker, 1986),
@@
Figure 6: Band structure (left) and corresponding longitudinal conductance (right, in units of $e^{2}/h$ ) for a zigzag edge nanoribbon with $\alpha=0.1$, $\beta=0.09$ in units of tata and the width is $N=21$. The brown dashed lines highlight interband anticrossings while the black dashed lines show interband spin degeneracy point.
By using Eq. (15) we compute zero temperature conductance as a function of chemical potential against the corresponding band structure for zigzag edge nanoribbon in the Fig. 6. As usual, the conductance exhibits quantization in units of $2e^{2}/h$ when $\mu<0.1$ . This can be attributed to the fact that only one subband with two electron incident channels is occupied by the chemical potential. As the chemical potential starts occupying the spin up subband of the lowest subband, both bands contribute one incident channel each maintaining the conductance again at $2e^{2}/h$. At $\mu\approx 0.1$ the conductance suddenly increases to $4e^{2}/h$ reflecting the existence of four electron incident channels, three from the spin up subband and one from the spin down subband. However, within very a very narrow range of $\mu$ around $0.1$, the conductance drops back to $2e^{2}/h$ and then rises up again to $4e^{2}/h$. This unusual feature originates from the presence of an anticrossing point at $\mu=0.1$. With further increase in $\mu$, the phenomenon reappears at $\mu=0.3$ corresponding to another anticrossing point. Such phenomenon occurs because of the drops of number of electron incident channel from the subband occupied by chemical potential. We emphasize that such behavior does not occur at the inter-subband spin degeneracy points. Hence we can conclude that although a signature of anticrossing point is visible in longitudinal conductance, it shows no corresponding features at the inter-band spin degeneracy. Finally, we note that the qualitative features of the conductance remain the same for the straight edge case, with the only difference being a shift in the locations of the anticrossing points.
@@ V Summary
We have investigated the spin-Hall transport and longitudinal transport for a 2D electronic system with RSOI and DSOI in a discrete lattice model. We have considered square lattice system with straight edge and zigzag edge geometry. In both cases, we showed that a number of spin degeneracy points and anticrossing points between two subbands with opposite spin emerges. The SHC shows a sharp enhancement when the chemical potential is situated at those spin degeneracy points or anticrossing points, exhibiting spin-Hall resonance. Hence, we conclude that resonance in SHC can be achieved without applying any external perturbation like magnetic field or light field as proposed previously. However such resonance can occur only if RSOI and DSOI both are present. We have also investigated the effects of anisotropy and finite temperature and demonstrated that the latter results in suppression of SHC but this suppression is weaker in higher energy degeneracy and anticrossing points while the former results in a shift of the position of resonance points in the SHC. We also study the spin Nernst effect and find that the spin Nernst coefficient reflects these anticrossing and spin degeneracy points through clear thermoelectric signatures. Finally we study the longitudinal conductance by using retarded Green function approach, and reveal that anticrossing point can give a signature to the conductance whereas the degeneracy point does not leave any mark.
@@ VI Acknowledgement
SK Firoz Islam acknowledges financial support by the project ANRF/ECRG/2024/005166/PMS.
Appendix A Straight edge nanoribbon: Band structure and Transport properties
We briefly discuss the case of a straight edge nanoribbon here. We present lowest few spin-split subbands for straight edge nanoribbon in Fig. 7 where red and blue curves represent the spin-up and spin-down components, respectively. Very much similar to the case of zigzag ribbon, here also a number of spin splitting vanishing as well as anticrossing points are present between two subbands, which are highlighted by circles and squares respectively. It is noticeable that the number of such points are relatively higher in straight edge case compared to zigzag edge. It is expected that the SHR can occur when the chemical potential passes through the inter subband spin degenerate or anticrossing points. The SHC is plotted in the Fig. 8 showing multiple resonances.
@@ Appendix A Straight edge nanoribbon: Band structure and Transport properties
Figure 7: spin resolved band dispersion of a straight edge nanoribbon for $\alpha=0.11$ and $\beta=0.1$ in units of tata. The color scale represents the expectation value of the in-plane spin operator, i.e., the spin polarization in units of $\hbar/2$. Anticrossing points are highlighted by squares, while spin degeneracy points are marked by circles.
@@
Figure 8: Spin resolved band structure (left) and SHC (in units of $\sigma_{0}=e/8\pi$) for a straight edge nanoribbon are plotted for width $N=21$, $\alpha=0.11$, $\beta=0.1$ in units of $ta$, at temperature $k_{B}T/t=10^{-5}$. The spin Hall resonances arising from anticrossings occur at chemical potentials $\mu/t=0.0418,0.14105,0.28511,$ and $0.47312$, highlighted by brown dashed lines. The resonance associated with vanishing spin splitting appears at $\mu/t=0.15919,0.359$ and $0.42526$, indicated by a black dashed line.
@@
Figure 9: spin resolved band structure (left) and corresponding longitudinal conductance (right, in units of $e^{2}/h$ ) for a straight edge nanoribbon with $\alpha=0.11$, $\beta=0.1$ in units of tata and the width is $N=21$. The brown dashed lines highlight interband anticrossings while the black dashed lines show interband spin degeneracy point.
Appendix B Full Band Structure of Zigzag- and Straight Edge Nanoribbons
In this appendix, we present the full spin resolved band structures of both zigzag edge and straight edge nanoribbons to provide an overall view of the energy spectrum and to illustrate the role of edge geometry. Figures 1012 display the complete band dispersions across the Brillouin zone. For the zigzag edge nanoribbon, edge localized states appear at $E=4t$, while such states are absent in the straight edge geometry. This distinction is consistent with earlier results reported for systems with pure RSOI in Ref. Kubo (2024), and we find that this qualitative feature remains unchanged in the simultaneous presence of RSOI and DSOI. Since these edge localized states occur at relatively high energies and accessing such large chemical potentials is not physical for typical semiconductor nanostructures, the main analysis in the paper is restricted to the lowest few subbands.
We further clarify the qualitative difference between the full band structures of straight edge and zigzag edge nanoribbons. Although the lowest few subbands in both geometries are qualitatively similar and exhibit conventional parabolic dispersion, a marked difference emerges at higher energies. In the zigzag edge nanoribbon, half of the subbands become inverted, above the flat edge band. Such band inversion is absent in the straight edge nanoribbon, where all subbands retain non-inverted, parabola like dispersion throughout the spectrum. As a consequence, within a given energy interval, the straight edge nanoribbon accommodates a larger number of subbands compared to the zigzag edge case. This increased subband density naturally leads to a higher number of band crossings and anticrossings, which is clearly reflected in Fig. 7.
When $\alpha\neq\beta$, pronounced anticrossings and spin degeneracy points emerge between spin split subbands accompanied by strong spin mixing, as indicated by the continuous variation of the spin polarization near the crossing and anticrossing points. In contrast, for $\alpha=\beta$, the band structure exhibits numerous spin degeneracy points distributed throughout the Brillouin zone. In this case, anticrossings are absent and only spin degenerate crossings appear. The degeneracies at any finite $k_{y}$ are of the same order as those at $k_{y}=0$ and persists for both zigzag and straight edge nanoribbon geometries. Moreover we see that when $\alpha=\beta$ the mixed spin states are absent near crossing points. As a consequence, the interband matrix elements responsible for SHC are suppressed, leading to a vanishing spin Hall response despite the presence of multiple spin degeneracy points in the band structure.
Figure 10: Spin resolved band dispersion of a zigzag edge nanoribbon is presented for $\alpha=0.1$ and $\beta=0.09$ in units of tata. The color scale represents the expectation value of the in-plane spin operator, i.e., the spin polarization in units of $\hbar/2$.
Figure 11: Spin resolved band dispersion of a zigzag edge nanoribbon is plotted for $\alpha=0.1$ and $\beta=0.1$ in units of tata. The color scale represents the expectation value of the in-plane spin operator, i.e., the spin polarization in units of $\hbar/2$.
Figure 12: Spin resolved band dispersion of a straight edge nanoribbon is plotted for $\alpha=0.11$ and $\beta=0.1$ in units of tata. The color scale represents the expectation value of the in-plane spin operator, i.e., the spin polarization in units of $\hbar/2$.
Figure 13: Spin resolved band dispersion of a straight edge nanoribbon is plotted for $\alpha=0.11$ and $\beta=0.11$ in units of tata. The color scale represents the expectation value of the in-plane spin operator, i.e., the spin polarization in units of $\hbar/2$.